OpenStudy (anonymous):

(3.8*10^8)*(6.9*10^4) Use scientific notation to multiply the following two numbers

OpenStudy (anonymous):

When you multiply scientific notation, simply multiply the numbers (3.8 and 6.9) and then add the exponents on the 10s (8+4).

OpenStudy (anonymous):

3.8 * 6.9 = 26.22 10^8 * 10^4 = 10 ^(8+4) = 10^12 So, 26.22*10^12

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

You're welcome. Do you feel like you understand how to do that now?

OpenStudy (anonymous):

I have a hard time understanding math and this forum helps me so much! Yes when you show it out it helps a lot!

OpenStudy (anonymous):

Great! If you have another similar problem, feel free to post it and work it out.

OpenStudy (anonymous):

Ok will do thanks!